OpenStudy (anonymous):

Factor completely: x2 − 36 (x − 6)(x − 6) (x − 6)(x + 6) (x − 36)(x + 1) (x − 36)(x − 1)

OpenStudy (anonymous):

THis is a difference in squares pattern. Just find the square root of 36 and put it where the question marks are. (x^2 - 36) = (x + )(x - ) The pattern is: (a^2 - b^2) = (a + b)(a - b)

OpenStudy (anonymous):

Or rather the blanks.

OpenStudy (anonymous):

i got the second one as my answer im i right

OpenStudy (anonymous):

Yup :)